OpenStudy (anonymous):

tell whether the given equation is a linear or not if it is change it to y=mx+b 1.x+y= 2.3x+2y-2=0

OpenStudy (anonymous):

neither if you have a "dangler" or b its not linear

OpenStudy (anonymous):

if you want #2 in the form of y=mx+b, simply isolate the y by moving the 2 two and 3x onto the other side of the "=" sign: 2y=2-3x (add 2 to both side and subtract 3x to both sides) y=(2-3x)/2 (divide by 2)

OpenStudy (anonymous):

\[y=-3(2x-4)= y=-6x+2 thats the example of a given problem of linear equation.. that's wht i need \]

OpenStudy (anonymous):

Um y=-3(2x-4) =/= y=-6x+2

OpenStudy (anonymous):

Well I'll assume it was a typo but if not, let it be known that y=-3(2x-4) = y=-6x+12